26.42

A resistor and capacitor are connected in series to an emfsource. The time constant for the circuit is 0.870 rm s.a-A second capacitor, identical to the first, is added inseries. What is the time constant for this new circuit?b-In the original circuit a second capacitor,identical to the first, is connected in parallel with the firstcapacitor. What is the time constant for this newcircuit?
Answer #1
We know the formula of the time constant for the RC circuitisT= RC = 0.870sa)When a capacitor is connected in series with another identicalcapacitor then the equivalent capacitance is Ceq =C / 2Then the effective time constant is T' =CeqR=(C/ 2)R=(1/2)RC=(1/2)T=(1/2)(0.870s) = 0.435sb)When a capacitor is connected in series with another identicalcapacitor then the equivalent capacitance is Ceq =2CThen the effective time constant is T' =CeqR=(2C)R=2RC=2T=(2)(0.870s)
answered by: Nakiya